# Creation of Wavepackets for Quantum Chromodynamics on Quantum Computers

### v1

@@ Abstract
One of the most ambitious goals of quantum simulation of quantum field theory is the description of scattering in real time, which would allow not only for computation of scattering amplitudes, but also for studying the collision process step by step. The initial state of such simulation is made of typically two wavepackets moving on top of the vacuum, whose preparation is hard. Some proposals are based on a modified adiabatic transformation of free wavepackets into interacting wavepackets, but this is not suitable for quantum chromodynamics (QCD). Other approaches used in one-dimensional lattice gauge theories rely on an adiabatic transformation from the strong coupling limit, or on ansatzes to create the wavepackets from the full, interacting vacuum. Here we extend a previous work to create wavepackets of a general kind of particle from the vacuum in lattice gauge theories in one and three dimensions, including QCD. The conceptual foundation of this approach is the Haag-Ruelle scattering theory, and the only theoretical limitation is given by massless particles. In the context of digital quantum computation, the wavepacket creation from the vacuum is implemented with the technique known as LCU (linear combination of unitaries). The preparation is performed successfully with a certain success probability, which may represent the main obstacle of our approach.
@@ I Introduction
Quantum computers are expected to have a large impact on the problems that present an exponential improvement with respect to classical computation. Of the few known problems, simulation of complex quantum systems is probably the most prominent. Among these, strongly-correlated many-body systems are very hard to deal with, because traditional approaches fail except in a handful of cases, or for limited classes of questions. Quantum and quantum-inspired approaches are promising solutions to this dilemma [1, 2, 3, 4, 5, 6, 7, 8].
@@
A quantum field theory is essentially a strongly-correlated quantum system with particular requirements on the symmetries it enjoys, namely Poincaré invariance. At a deep level, the correct implementation of this symmetry group also seems to demand gauge invariance to obtain nontrivial dynamics, at least in a three-dimensional space. The most challenging quantum field theory is quantum chromodynamics (QCD), which describes the strong interaction between quarks and gluons. Differently from all the other sectors of the Standard Model of particle physics, a large part of its phenomenology is intrinsically nonperturbative, whence analytical techniques can be used only to set up the right framework for numerical calculations. The gauge group is $SU(3)$, which increases computational costs by several orders of magnitude with respect to a $U(1)$ theory, to make a comparison. Traditional lattice calculations can only address problems amenable to the imaginary-time formalism, such as prediction on particle masses and matrix elements. Nonperturbative studies of processes that depend on real-time evolution is one of the major open problems of particle physics.
@@
For the reasons just mentioned, considerable effort within the quantum-computing community and the high-energy-physics one points towards the ultimate goal of simulating real-time evolution of processes within QCD [9, 10, 11, 12, 13]. The resources required to perform a useful quantum calculation in QCD are several orders of magnitude beyond the present capabilities [14], but promising progress has been made in the last couple of years towards the realization of a large-scale, error-corrected quantum computer, both on the experimental and the theoretical side. Although a full quantum computer is still out of reach, it nonetheless makes sense to analyze all the possibilities to obtain realistic resource estimates, and determine which questions we will be able to address.
Specifically, we consider scattering of particles, a class of experiments which is extremely important for our understanding of fundamental physics. At very high energy, QCD is perturbative, and perturbation theory is well suited to study cross sections and test the theory against experimental results. At low energies, perturbation theory is no longer useful, so it is very challenging to study, for example, pion-pion or proton-proton collisions above the inelastic-scattering threshold. Quantum simulation offers a possibility to study the real-time evolution of these processes, but compared to other applications, the initial state preparation is particularly difficult. In the framework of the Hamiltonian formulation on a spatial lattice, our goal is to prepare two wavepackets of some kind of particles on top of the vacuum. For other approaches that avoid the necessity of wavepackets, see [15, 16]. In most existing approaches, the key step consists of a modified adiabatic transformation to turn wavepackets of the free theory into wavepackets of the interacting theory [17, 18, 19, 20, 21]. Since in the free theory there are no states corresponding to composite particles, this method can only be used for particles associated with elementary degrees of freedom. As a consequence of confinement in QCD, no quarks or gluons appear in the spectrum, and all the particles (hadrons) are composite. Therefore, the works in [17, 18, 19, 20, 21] do not apply to QCD.
There are a few ideas in the literature to circumvent this problem: variational approaches [22], variational approaches combined with adiabatic transformation from the strong-coupling limit [23], and optimized ansatzes [24]. An approach which does not require optimization or ansatzes is [25], based on the Haag-Ruelle scattering theory. In all these methods, wavepackets are prepared directly from the vacuum of the full, interacting theory, which avoids the necessity for adiabatic transformations. In the context of theories with a mass gap, the approach in [25] is fully general, and easily adaptable to different formulations and different particles. Furthermore, its asymptotic efficiency can be analyzed theoretically. On the other hand, the works in [22, 23, 24] have been tested numerically only in simple models in one dimension so far, and it is difficult to perform the analysis of the asymptotical complexity. Generalizations to higher dimensions are possible but not straightforward.
In this work, we show how to apply the method in [25] to lattice gauge theories. As a warm up, we consider $U(1)$, $SU(2)$ and $SU(3)$ theories in one dimension. Subsequently, we address QCD in three dimensions with two flavors of quarks. Not only the one-dimensional theories provide simpler versions of QCD where to make some tests, but also the tools that are used in the one-dimensional case are also used to build up the techniques in the three-dimensional case. As examples, we consider the pions and the nucleons, while other kinds of particles can be treated in a completely analogous way.
We work in the Kogut-Susskind formulation, since this is the choice of preference in many works on quantum simulation of lattice gauge theories in the Hamiltonian formulation. However, our conclusion is that in three dimensions this is not a good choice, because the lattice fermionic degrees of freedom do not correspond with the convenient degrees of freedom in the continuum (Dirac spinors). In addition to the problems already addressed in [26], the Kogut-Susskind formulation has two more disadvantages when we deal with particle scattering: 1) large overheads in the creation of wavepackets with our approach; 2) mixing of flavor symmetries and cubic symmetries, which makes the concepts of spin and isospin even less precise. While the first problem is specific to our method, the second one is not. Applying our ideas to formulations other than the Kogut-Susskind one, like Wilson fermions, is simpler.
@@
The rest of this paper is organized as follows. In section II, we review the approach presented in [25], and we extend it to particles of general spin. In section III we give the lattice operators that are necessary for the creation of wavepackets in one-dimensional lattice gauge theories $U(1)$, $SU(2)$ and $SU(3)$. In section IV we show how to implement the same operators in terms of qubit operators. Finally, in section V we consider QCD in three dimensions. In the appendix we review the generalized superfast encoding of fermions onto qubits, and we also show how to implement odd operators.
@@ II Creation of Wavepackets from the Vacuum
In this section we briefly review the main features of the method proposed in [25], based on the Haag-Ruelle scattering theory, to create wavepackets of particles in an interacting field theory from the vacuum of the theory. The Haag-Ruelle formulation [27, 28] was originally developed in axiomatic quantum field theory and later on formulated for lattice systems [29]. For a review of the formulation in the continuum, see [30].
We can create wavepackets of a certain particle using an operator of the form
@@
where $D=d+1$ is the number of spacetime dimensions, $x$ is a spacetime coordinate, $\psi$ is a Schwartz function, and $\hat{\mathcal{O}}$ is an interpolating Heisenberg operator for the particle we want to create. The function $\psi$ and the operator $\hat{\mathcal{O}}$ are chosen considering the spectrum and the symmetries of the theory.
For example, suppose there is only one particle of mass $m$ in the spectrum. Then, the joint energy-momentum spectrum is made of three disjoint sectors:
@@
  • the multi-particle continuum, defined by $p_{\mu}p^{\mu}\geq 4m^{2}$.
The function $\psi$ is chosen as the Fourier transform of a function $\tilde{\psi}$ with support overlapping only with the one-particle mass-hyperboloid, and not with the multi-particle continuum, as depicted in figure 1. The interpolating operator $\hat{\mathcal{O}}$ should carry the same quantum numbers as the particle. More precisely, if $|\alpha\rangle$ is a one-particle state and $|\Omega\rangle$ the vacuum, then the condition for $\hat{\mathcal{O}}$ to be an interpolating operator is
@@
Figure 1: Structure of the joint spectrum of $P_{\mu}$ in a one-dimensional space. The blue line represents the one-particle mass hyperboloid, the red region is the multiparticle continuum, and the green spot represents the support of the function $\tilde{\psi}$.
When we pass from the continuum spacetime to the lattice, we use the same approach. The main difficulty is that on the lattice we typically have only discrete versions of the symmetries in the continuum, and there may be different kinds of particles that are not present in the continuum limit. When we discretize the theory on a space lattice with spacing $a$ to make a quantum simulation, equation (1) becomes
[…]
where $\hat{\mathcal{O}}(\mathbf{x})$ is the Schrödinger representation of $\hat{\mathcal{O}}(x)$, and $x=(t,\mathbf{x})$. The lattice theory needs to be mapped onto a qubit system, with all due truncation and digitization [31, 32]. Also the time integral in equation (4) needs to be approximated by a discrete and truncated summation in steps of $\delta_{t}$,
$\hat{a}_{\psi}^{\dagger}=\sum_{\mathbf{x},t}a^{d}\delta_{t}\psi(t,\mathbf{x})e^{iHt}\hat{\mathcal{O}}(\mathbf{x})e^{-iHt}.$ (5)
Once we know how to implement $\hat{\mathcal{O}}$ on a quantum computer, the creation operator $\hat{a}_{\psi}^{\dagger}$ can be implemented by means of a linear combination of unitaries (LCU) [33]. If $T$ is the size of $\psi$ in the time direction, this requires simulating the evolution for a time $O(T)$.
The interpolator $\hat{\mathcal{O}}(\mathbf{x})$ can be typically written as a LCU itself,
$\hat{\mathcal{O}}(\mathbf{x})=\sum_{j}c_{j}\mathcal{U}_{j}(\mathbf{x}),\qquad\sum_{j}|c_{j}|=C.$ (6)
Overall, $\hat{a}_{\psi}^{\dagger}$ is successfully implemented with probability
@@
where $\delta_{p}$ is the extension of the support of $\tilde{\psi}$ in the space directions, $\bar{E}$ is the energy of the particle, and $Z$ is the squared renormalization constant connecting $\hat{\mathcal{O}}(\mathbf{x})$ with its continuum normalization, typically logarithmic in the lattice spacing [34].
Generally, $\hat{\mathcal{O}}(x)$ can be a spinorial field (or another kind of field), as is the case for the proton, and therefore the result needs to be adapted. Assume that under a Lorentz transformation in the continuum we have $\hat{\mathcal{O}}_{\alpha}\to\sum_{\beta}S(\Lambda)_{\alpha\beta}\hat{\mathcal{O}}_{\beta}$, and that the operators $\hat{\mathcal{O}}_{\alpha}$ interpolate for a particle of spin $J$ and mass $m$. There are $12J+$ 1 one-particle states with momentum $\mathbf{k}$, $|\mathbf{k},j\rangle$, with $j=-J,-J+1,\dots,J$. Then we have
@@
We consider now the example of a spinorial field in the Dirac representation interpolating for a spin $1/2$ particle with momentum $\bar{\mathbf{k}}=(k_{x},0,0)$. We have
@@
We see that, for a spinorial interpolator, and if $Z_{1}^{j}\neq Z_{4}^{j}$, the term $k_{x}/m$ from this expansion compensates the term $\bar{E}$ in the numerator of equation (8) for highly energetic spin-$1/2$ particles.

### v2

@@ Abstract
One of the most ambitious goals of quantum simulation of quantum field theory is the description of scattering in real time, which would allow not only for computation of scattering amplitudes, but also for studying the collision process step by step. The initial state of such simulation is made of typically two wave packets of stable particles moving on top of the vacuum, whose preparation is difficult. Here we extend a previous work to create wave packets of a general kind of particle from the vacuum in lattice gauge theories in various dimensions, including three-dimensional QCD for the first time. The conceptual foundation of this approach is the Haag-Ruelle scattering theory, and the only theoretical limitation is given by the presence of massless particles. In the context of digital quantum computation, the wave packet creation from the vacuum is implemented with the technique known as LCU (linear combination of unitaries). The preparation is performed successfully upon measuring an ancillary register with a certain probability, which vanishes polynomially in the lattice spacing, the wave-packet energy and the momentum narrowness.
@@ I Introduction
Quantum computation is a new promising paradigm to render feasible some computational problems that are classically too hard to be solved even with the huge amount of resources of the best supercomputers. More precisely, quantum computers are expected to have a truly strong impact in problems that require an exponentially large amount of classical resources and a polynomially large amount of quantum resources. For these problems, having a proper quantum computer would make the difference between solving relevant cases and not solving them. It would not be just a matter of making the computation faster. However, instances of such problems are rare. Of the few known examples, simulation of complex quantum systems, especially of strongly correlated many-body systems, is probably the most prominent. Traditional approaches, numerical or analytical, fail except in a handful of cases, or for limited classes of questions. For these reasons, quantum and quantum-inspired approaches are being widely studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
A quantum field theory is essentially a strongly correlated quantum system with particular requirements on the symmetries it enjoys, namely Poincaré invariance. At a deep level, the correct implementation of this symmetry group also seems to demand gauge invariance to obtain nontrivial dynamics, at least in a three-dimensional space. A particularly challenging case is quantum chromodynamics (QCD), which describes the strong interaction between quarks and gluons. Differently from all the other sectors of the Standard Model of particle physics, a large part of its phenomenology is intrinsically nonperturbative, whence analytical techniques can be used only to set up the right framework for numerical calculations. The gauge group is $SU(3)$, which increases computational costs by several orders of magnitude with respect, for instance, to a $U(1)$ theory. Traditional lattice calculations are well suited to address problems in the imaginary-time formalism, such as prediction on particle masses and matrix elements, and information on the real-time evolution can be extracted in some cases [15]. However, general nonperturbative studies of processes in real-time evolution is one of the major open problems of particle physics.
@@
For the reasons just mentioned, considerable effort within the quantum-computing and the high-energy-physics communities aims at simulating real-time evolution of processes within QCD [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. The resources required to perform a useful quantum calculation in QCD are several orders of magnitude beyond the present capabilities [34], but promising progress has been made in the last years towards the realization of a large-scale, error-corrected quantum computer, both on the experimental and the theoretical side [35, 36]. Although a full quantum computer is still out of reach, it nonetheless makes sense to analyze all the possibilities to obtain realistic resource estimates, and determine which questions we will be able to address. Furthermore, on the long way to address full QCD there are many simpler theories, such as the Schwinger model, that can be used to test and benchmark new techniques, and to gain insight on some aspects of QCD and quantum field theory in general.
Specifically, we consider scattering of particles, a class of experiments which is extremely important for our understanding of fundamental physics. At very high energy, QCD is perturbative. The Lehmann-Symanzik-Zimmermann (LSZ) reduction formula [37] is a powerful tool to study cross sections at asymptotic times in perturbation theory and to obtain predictions to be tested against experimental results. However, at low energies, perturbation theory is no longer useful, so it is very challenging to study, for example, pion-pion or proton-proton collisions above the inelastic-scattering threshold. The Haag-Ruelle formalism [38, 39, 40] is another excellent framework in which to study scattering. It was developed in the context of axiomatic quantum field theory, and it provides the logical link between the Wightman axioms and the LSZ reduction formula. While being a conceptual milestone of quantum field theory, it is not well suited for perturbative calculations, this being the main reason why it was outshone by the LSZ formalism. However, quantum computation may be the missing element to exploit the Haag-Ruelle theory in a practical way too, as we will shortly see.
Quantum simulation offers a possibility to study the real-time evolution of scattering processes, but compared to other applications, the initial state preparation is particularly difficult. In the framework of the Hamiltonian formulation on a spatial lattice, our goal is to prepare two wave packets of stable particles on top of the vacuum. If the goal is just to compute scattering amplitudes, wave packet preparation can be avoided [41, 42]. However, simulation of wave-packet collisions would allow for much more than computation of scattering amplitudes. We could directly monitor the dynamics step by step, and gain insight on what happens during the collision, which is not easily accessible even in perturbative regimes.
Concerning the preparation of wave packets, the key step in the first proposed approaches consists of a modified adiabatic transformation to turn wave packets of the free theory into wave packets of the interacting theory [43, 44, 45, 46, 47]. Since in the free theory there are no states corresponding to composite particles, this method can only be used for particles associated with elementary degrees of freedom. As a consequence of confinement in QCD, no quarks or gluons appear in the spectrum, and all the particles (hadrons) are composite. Therefore, the approach used in [43, 44, 45, 46, 47] does not apply to QCD.
To circumvent the problem, other strategies have to be followed. Several ideas have been discussed and explored numerically in one dimension. In [48, 49], variational algorithms are used to obtain wave packets in a scalar-field theory. Variational approaches are also used in the Schwinger model in [50]. Preparation of $W$ states combined with variational techniques is used in [51], which works well in models with a short correlation length and, thus, far from the continuum limit. In the context of the Abelian gauge theories $Z_{2}$ and $U(1)$, ideas based on ansatz-optimization [52, 53] or quantum subspace expansion [54] have been proposed to obtain operators that create mesonic wave packets from the interacting vacuum, similarly to what is done in [2] using tensor networks and in [55] for the Thirring model. Generalization of these methods to more complicated cases, such as mesonic and baryonic particles in multidimensional non-Abelian theories, is possible in principle, but it has to be showed that these methods remain efficient when more complicated structures are considered, and when larger wave packets in coordinate space are needed to obtain more energetic wave packets with a narrower distribution in momentum space in the continuum limit. In [50], another idea is proposed for theories displaying confinement: wave packets are adiabatically transformed from the strong-coupling limit to the full theory, rather than from the free theory. This should avoid the issue explained in the previous paragraph, but more investigation is required. A possible problem is that in the continuum limit a theory like QCD is weakly coupled due to asymptotic freedom, and the adiabatic transformation may be difficult to implement.
A general approach, based on the Haag-Ruelle scattering theory, has been introduced in our previous work [56], where we described how to obtain creation operators for elementary and composite particles in interacting scalar-field theories. These operators act on the interacting vacuum, so efficient interacting-vacuum preparation is assumed to be available. Another required ingredient is knowledge of the spectrum. The only theoretical limitation of this strategy is the existence of an isolated mass shell, which should be guaranteed by a finite mass gap in the continuum limit. Apart from this, it is easily adaptable to different theories, different formulations, and different particles, as we show in this work. The creation operators are obtained by appropriately smearing interpolating operators for a given particle over space and time. Its asymptotic efficiency can be shown in the continuum limit and for high-energy wave packets with narrow distribution in momentum space. The main difficulty is that it only succeeds upon measuring an ancillary register of qubits with a certain outcome. The success probability is polynomially vanishing in the lattice spacing, the wave-packet energy and the momentum narrowness.
In the present work, we show how to apply the method based on the Haag-Ruelle formalism [56] to QCD. We provide the first description of an efficient strategy to prepare wave packets in this theory, assuming there is an efficient way to prepare the vacuum state. Apart from addressing theories different from our previous work [56], the new aspects we address here concern how to deal with particles with spin one-half and higher, how to write interpolating operators in terms of qubit operators, and estimating the success probability for different particles. To achieve this, we bring together tools and concepts from different areas of research developed over several decades, ranging from axiomatic quantum field theory, traditional lattice field theory and Hamiltonian formulation of lattice theories to quantum computation. As a warm up, we consider $U(1)$, $SU(2)$ and $SU(3)$ theories in one dimension. We use them to introduce some building blocks required for three-dimensional QCD. Moreover, the simulation of these theories requires much less resources than full QCD, and they are often used as toy models to test ideas. Treating them separately from QCD may be useful for implementations in the near or mid-term future. We focus, for simplicity, on the well-known pions and nucleons. Other kinds of particles can be treated in a completely analogous way.
It is not clear which formulation of lattice QCD will be best suited for quantum simulation, but a common choice has been the Kogut-Susskind formulation, a choice we henceforth follow. Other possibilities are the loop-string-hadron formulation [57, 58, 59, 60], or the Wilson-fermion formulation [61]. The former is based on the Kogut-Susskind formulation, so developing tools for the Kogut-Susskind formulation is typically necessary or useful also for the loop-string-hadron one. Porting the tools developed in this paper to the Wilson-fermion formulation is straightforward. Another option is to use the orbifold formulation [62, 63], which can also be connected to the Kogut-Susskind framework. Therefore, the choice of the Kogut-Susskind formulation is actually quite general.
@@
The rest of this paper is organized as follows. In section II, we review the approach presented in [56], and we extend it to particles of general spin. In section III we give the lattice operators that are necessary for the creation of wave packets in one-dimensional lattice gauge theories $U(1)$, $SU(2)$ and $SU(3)$. In section IV we show how to implement the same operators in terms of qubit operators. Finally, in section V we consider two-flavoured QCD in three dimensions, and in section VI we give our conclusions. In the appendix we review the generalized superfast encoding of fermions onto qubits, and we also show how to implement odd operators.
@@ II Creation of Wave Packets from the Vacuum
In this section we briefly review the main features of the method proposed in [56], based on the Haag-Ruelle scattering theory, to create wave packets of particles in an interacting field theory from the vacuum of the theory. In  [56] scalar-field theories were considered, while in this paper we take at first a more general view. Then, at the end of the section, we focus on spin-one-half particles, and give some results that are necessary to determine the complexity of the method. The Haag-Ruelle formulation [38, 39] was originally developed in axiomatic quantum field theory and recently formalized for lattice systems [64]. For a thorough review of the formulation in the continuum see [40], whose extensive explanation goes perhaps beyond the purposes of quantum simulation. We invite the reader interested in quantum simulation to go through [56], which also includes a more concise review of the Haag-Ruelle formalism.
For our purposes, the relevant result is that we can create wave packets of a certain particle using operators of the form
@@
where $D=d+1$ is the number of spacetime dimensions, $x$ is a spacetime coordinate, $\psi$ is a Schwartz function, and $\hat{\mathcal{O}}$ is an interpolating Heisenberg operator for the particle we want to create. The function $\psi$ and the operator $\hat{\mathcal{O}}$ have to be properly chosen considering the spectrum and the symmetries of the theory. The operator in equation (1) is sometime called a Haag-Ruelle creation operator. For proper choices of $\psi$ and $\hat{\mathcal{O}}$, the state $\hat{a}_{\psi}^{\dagger}|\Omega\rangle$, where $|\Omega\rangle$ denotes the vacuum, is a one-particle state. To create multiple wave packets in different space regions, we just need to apply multiple Haag-Ruelle creation operators with spatially separated functions $\psi_{1},\,\psi_{2},\dots$ Intuitively, the function $\psi$ is responsible for selecting one-particle states and cutting off multiparticle states. The operator $\hat{\mathcal{O}}$ is responsible for selecting the desired kind of particle. Some examples should help to understand these points.
To begin, suppose we have a theory with only one particle of mass $m$ in the spectrum. The momentum operators $P_{0}=H$ and $P_{i}$, $i=1,\dots,d$, commute with each other, so we can consider their joint spectrum. This is made of three separate sectors:
@@
  • the multiparticle continuum, defined by $p_{\mu}p^{\mu}\geq 4m^{2}$.
The function $\psi$ is chosen as the Fourier transform of a function $\tilde{\psi}$ with support overlapping only with the one-particle mass hyperboloid, and not with the multiparticle continuum, as depicted in figure LABEL:joint-spectrum.
Figure 1: Two examples of the joint energy-momentum spectrum in continuum quantum field theory. The blue lines represent the one-particle mass hyperboloids, the red regions are the multiparticle continua, and the green spot represents the support of the function $\tilde{\psi}$. For visual clarity, only one spatial component is shown.
The interpolating operator $\hat{\mathcal{O}}$ should carry the same quantum numbers as the particle. More precisely, if $|\alpha\rangle$ is a one-particle state and $|\Omega\rangle$ the vacuum, then the condition for $\hat{\mathcal{O}}$ to be an interpolating operator is
@@
Finally, we can consider QCD with two flavours of quarks. The low-lying spectrum is depicted in figure LABEL:QCD-spectrum, where only one spatial component is shown for visual clarity. The fact that the one-nucleon mass hyperboloid lies in the multipion continuum is not a problem because the two kind of particles have different spins and isospins.
On a lattice, the same ideas apply with due modifications. Coordinate space and momentum space are discretized, while time is left continuous. Spectra are distorted, but should still contain separate one-particle and multiparticle sectors. We may then rewrite equation (1) on a lattice with spacing $a$ as
$\hat{a}_{\psi}^{\dagger}=\sum_{\mathbf{x}}a^{d}\int_{-\infty}^{+\infty}\!dt\,\psi(t,\mathbf{x})e^{iHt}\hat{\mathcal{O}}(\mathbf{x})e^{-iHt},$ (4)
where $\hat{\mathcal{O}}(\mathbf{x})$ is the Schrödinger representation of $\hat{\mathcal{O}}(x)$, and $x=(t,\mathbf{x})$. When moving from the continuum to the lattice, many symmetries are lost, and new particles may appear. As a consequence, there arises the problem of identifying particles and using the correct lattice interpolators. We deal with this more in detail in section V.
The lattice theory has to be mapped onto a qubit system, with all due truncation and digitization [65, 66]. Also the time integral in equation (4) has to be approximated by a discrete and truncated summation in steps of $\delta_{t}$ over $N$ points,
$\hat{a}_{\psi}^{\dagger}=\sum_{\mathbf{x}}\sum_{i=1}^{N}a^{d}\delta_{t}\psi(t_{i},\mathbf{x})e^{iHt_{i}}\hat{\mathcal{O}}(\mathbf{x})e^{-iHt_{i}}.$ (5)
The required number of time samples can be shown to grow like
$N=O\bigg{(}\frac{T}{\sqrt{\epsilon}}\sqrt{\big{\|}\big{[}H,[H,\hat{\mathcal{O}}(\mathbf{x})]\big{]}\big{\|}}\bigg{)},$ (6)
where $T$ is the size of $\psi$ in the time direction, and $\epsilon$ is the error introduced by the time-integral discretization. On a quantum computer, we can only implement operators with $\|\hat{\mathcal{O}}\|=O(1)$. Assuming the Hamiltonian is a sum of local terms, $H=\sum_{\mathbf{x}}H(\mathbf{x})$, we conclude that
$N=O\bigg{(}\frac{T}{\sqrt{\epsilon}}\|H(\mathbf{x})\|\bigg{)}.$ (7)
Notice that $N$ does not depend on the lattice size.
Once we know how to implement $\hat{\mathcal{O}}$ on a quantum computer, the creation operator $\hat{a}_{\psi}^{\dagger}$ can be implemented by means of a linear combination of unitaries (LCU) [67]. If $S$ is the number of lattice sites in the support of $\psi$, this requires an ancillary register of $\log_{2}(NS)$ qubits, $NS$ calls to controlled versions of $\hat{\mathcal{O}}$, and time evolution for a time $O(T)$. The time evolution is likely the most expensive part in the implementation of $\hat{a}_{\psi}^{\dagger}$, since $N$ depends only on the local terms of the Hamiltonian, while the time evolution depends on the whole Hamiltonian. Also, the local operators $\hat{\mathcal{O}}$ we consider in this work are not hard to implement.
The interpolator $\hat{\mathcal{O}}(\mathbf{x})$ is typically not unitary, so it has to be block encoded. Since any operator on a qubit system can be expanded as a linear combination of unitary operators, in general we can use LCU,
$\hat{\mathcal{O}}(\mathbf{x})=\sum_{j=1}^{L}c_{j}\mathcal{U}_{j}(\mathbf{x}),\qquad\sum_{j}|c_{j}|=C,\qquad c_{j}\in\mathbb{C}.$ (8)
In the cases we consider below, $L$ is a constant depending only on the kind of particle and the theory. Another $\log_{2}L$ ancillary qubits are required to implement $\hat{\mathcal{O}}(\mathbf{x})$. The quantity $C$ depends on $L$ and on the mass dimension of the interpolator.
Overall, $\hat{a}_{\psi}^{\dagger}$ is successfully implemented upon measurement of the ancillary register with probability
@@
The success probability determines the efficiency of our method for wave-packet preparation. If we want to create two wave packets, with success probabilities $\rho_{1}$ and $\rho_{2}$, we need to repeat the vacuum preparation and wave-packet creation about $1/(\rho_{1}\rho_{2})$ times before succeeding. This can be quadratically improved applying amplitude amplification [68].
@@
where $\delta_{p}$ is the linear size of the support of $\tilde{\psi}$ in the spatial components of the momentum, $\bar{E}$ is the energy of the particle, and $Z$ is the squared renormalization constant connecting $\hat{\mathcal{O}}(\mathbf{x})$ with its continuum normalization, typically logarithmic in the lattice spacing [69].
In general, $\hat{\mathcal{O}}(x)$ can be a spinorial field (or another kind of field), as is the case for the proton, and the result has to be adapted. Assume that under a Lorentz transformation in the continuum we have $\hat{\mathcal{O}}_{\alpha}\to\sum_{\beta}S(\Lambda)_{\alpha\beta}\hat{\mathcal{O}}_{\beta}$, and that the operators $\hat{\mathcal{O}}_{\alpha}$ interpolate for a particle of spin $J$ and mass $m$. There are $2J+1$ one-particle states with momentum $\mathbf{k}$, $|\mathbf{k},j\rangle$, with $j=-J,-J+1,\dots,J$. Then we have
@@
We consider now the example of a spinorial field in the Dirac representation interpolating for a spin-one-half particle with momentum $\bar{\mathbf{k}}=(k_{x},0,0)$. We have
@@
Notice that, for a spinorial interpolator, and if $Z_{1}^{j}\neq Z_{4}^{j}$, the term $k_{x}/m$ from this expansion compensates the term $\bar{E}$ in the numerator of equation (10) for highly energetic spin-one-half particles. If instead $Z_{1}^{j}=Z_{4}^{j}$, then we need to consider the second term in the expansion, which brings another suppression factor of $1/\bar{E}$ in the success probability.
In this section we have discussed the general features of how to prepare wave packets on quantum computers using the Haag-Ruelle scattering theory, focusing on particles of spin zero and one half. To continue the discussion, we need to consider specific examples. In particular, we need to have explicit expressions for the interpolating operators to estimate the quantity $C$ in (9). This is the content of the following sections.